OpenStudy (anonymous):

How do I partially differentiate f = 2x/(x^2 +y^2) wrt x and y? Is it by the quotient rule? Any help is very much appreciated.

OpenStudy (experimentx):

by treating one as constant

OpenStudy (anonymous):

quotient rule needed for \(f_x\) but not for \(f_y\) since in that case you view \(2x\) as a constant

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

who nellie, you are differentiating not integrating.

OpenStudy (anonymous):

\[f_x=\frac{(x^2+y^2)2-x^2(2x)}{(x^2+y^2)^2}\] then some algebra to clean it up

OpenStudy (anonymous):

thnxxxxxx

OpenStudy (anonymous):

\(f_y\) is easier because you treat the numerator as a constant you get \[f_y=-\frac{2y}{(x^2+y^2)^2}\]

OpenStudy (anonymous):

I ended up getting 2/(x^2 + y^2) - 4x^2/(x^2 +y^2)^2 for the partial derivative wrt x and -4yx/(x^2 +y^2) for y. Any idea where I might be going wrong? Thanks again

OpenStudy (anonymous):

maybe i made a mistake. lets do it for \(f_y\)

OpenStudy (anonymous):

oh right, doh you are rigth about \(f_y\)

OpenStudy (anonymous):

well, except the denominator should be squared

OpenStudy (anonymous):

Why is that?

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

or in this case just \[\frac{d}{dy}\frac{1}{f(y)}=-\frac{f'(y)}{f^2(y)}\]

OpenStudy (anonymous):

first one it looks like you are using the product rule, so i might be the same as my answer

OpenStudy (anonymous):

oh yes, you are right. i missed a 2 from the 2x up top you are correct

OpenStudy (anonymous):

Ah, fantastic. Thanks a million.